OpenStudy (anonymous):

If the point (-1/2, y) lies on the line whose equation is 2x - 3y = 6, what is the value of y? A.-7/3 B.-5/3 C.7/3

OpenStudy (anonymous):

A

OpenStudy (anonymous):

@mekaprasantho , don't give the answer! just show him\her the way!

OpenStudy (anonymous):

@DjNck , just put x = -1/2 in the equation!

OpenStudy (anonymous):

2(-1/2) - 3y = 6 solve for y :)

OpenStudy (anonymous):

2x - 3y = 6 Since (-1/2, y) lies on the line therefore substituting x= =-1/2 and y=y in the given eq we find 2*(-1/2) - 3y = 6 i.e. -1 -3y=6 i.e. -3y = 6+1 i.e. -3y =7 i.e. y=7/(-3) i.e. y=-7/3 Thus option A is correct

OpenStudy (anonymous):

Thanks everyone

OpenStudy (anonymous):

you're welcome :)

OpenStudy (anonymous):

kkPFEH199 i'll do that from next time

OpenStudy (anonymous):

thanks for advice

OpenStudy (anonymous):

you're welcome :) i hope i didn't upset you :)

OpenStudy (anonymous):

no

OpenStudy (anonymous):

im new to this site

OpenStudy (anonymous):

good :)